OpenStudy (anonymous):

what is the solution to the equation? 6y+1=-11

OpenStudy (anonymous):

\[6y+1=-11\] \[6y=-11-1\] \[6y=-12\] \[y=\frac{-12}6\] \[y=-2\]

OpenStudy (anonymous):

y= -11-1/6 y= -12/6 y= -2

OpenStudy (anonymous):

6y+1-1=-11-1 6y=-12 6y/6=-12/6 y=-2

OpenStudy (anonymous):

6y + 1 = -11 6y +1 -1 = -11 -1 6y = -12 6y/6 = -12/6 y = -2